If two equal charges, each of 1 C are separated in the air by a distance of 1 km, what would be the magnitude and kind of force between them?


1
Expert's answer
2022-03-06T15:17:56-0500

The foce between an equal charges are repulsive. According to the Coulomb's law:


"F = k\dfrac{q_1q_2}{r^2}"

where "q_1 = q_2 = 1C" are the charges, "k = 9\times10^9N\cdot m^2/C^2" is the constant, and "r = 1000m" is the distance between the charges. Thus, obtain:


"F = 9\times 10^9\cdot \dfrac{1\cdot 1}{1000^2}=9000N"

Answer. 9000N.